# Quadratic Curvature Correction and Its Breakdown to Thermodynamics of Rotating Black Holes

### v1

@@ 5 Corrections with fixed mass and angular momenta
It was pointed out that for the system with fixed mass, angular momenta and charges, one expects $\Delta S>0$ for stable thermodynamics [19]. Although our $\Delta S$ can be both positive and negative for given $\alpha$, this does not provide a counterexample of [19], since as we have pointed it out earlier that Myers-Perry black holes have unstable thermodynamics for all allowed parameters. However, a counterexample to [19] does exist, found in the higher-derivative correction in supergravity [26].
@@ 7 Divergence is a perturbative artefact: an explicit example
While we do not have exact solutions of rotating black holes in Einstein gravity extended with quadratic or more general higher-order derivative terms, we expect that the divergence we explored in this paper is an artefact of perturbation, indicating a breakdown only of the perturbative approach The divergence may not necessarily exist in the full nonlinear level.
@@ 8 Conclusions
In this paper, we began a review of Myers-Perry rotating black holes in diverse dimensions. We presented a simple analytic proof that the rotating black holes in even dimensions are all thermodynamically unstable. Specifically, we showed that the trace of the corresponding thermodynamic metric has an elegant expression (21) and it is manifestly non-positive. However, for odd dimensions, we have to employ numerical analysis to prove the instability. We also found that in odd dimensions when one and only one angular momentum vanishes, there existed an intriguing decoupling limit of extremal solutions whose geometry (16) has an unusual AdS3 embedding.
@@
In addition to the divergence of the $\alpha$-order correction, there are points in the parameter space that the correction vanishes. What is intriguing is that these zeros occur at the same points for the corrections to different thermodynamic variables in different ensembles. For example, $\Delta J$ in grand canonical ensemble vanishes at the exact same point as $\Delta\Omega$ in canonical ensemble. This can be seen from the numerator of $\Delta J$ in eqns (3.2.1), (3.2.1), (41), (3.3), and compare them to the corresponding $\Delta\Omega$ in eqns (70), (74), (79), (86). The similar relation exists between $\Delta M$ in canonical ensemble and $\Delta T$ in system with fixed mass and angular momenta, as we can see from eqns (70), (74), (79), (86) and (5), (5), (5), (5). These phenomena can be understood from the identities of differential equations, namely

### v2

@@ 2.4 Thermodynamic instability
As we shall demonstrate in section 5, the specific heat capacity can be positive for low temperature including the extremal limit. Therefore, the thermodynamic can be stable if we fix all the angular momenta to some specific values so that they do not involve in the thermodynamic process.
@@ 5.1 Corrections and zero-temperature divergence
5.2 On the sign of $\Delta S$
It was pointed out that for the system with fixed mass, angular momenta and charges, one should expect $\Delta S>0$ for stable thermodynamics [19]. We have already pointed it out earlier that Myers-Perry black holes have unstable thermodynamics for all allowed parameters, if we consider mass and all angular momenta as thermodynamic variables. In black hole thermodynamics, stability is rather rare. Even for the Reissner-Nordström (RN) black hole, its thermodynamics cannot be stable if both mass and charge are treated as thermodynamic variables, since the product of the specific heat capacity and charge capacitance is always negative [13]. One way to resolve this instability issue for the RN case is to fix the charge to a specific value so that it does not involve in the thermodynamic process, in which case, the specific heat capacity with fixed $Q$ is positive for low temperature. We can follow the same strategy, and fix all the angular momenta to some appropriate specific values to extract them from the thermodynamic process. As an example, for simplicity, we consider the case with all equal angular momenta $J_{i}=J/N$. For odd dimensions with $D=2N+1$, we find that the specific heat capacity is
$C_{J}={\frac{\partial M_{0}}{\partial T_{0}}}\Big{|}_{J}=\frac{\pi(D-2)\Sigma_{D-2}\left(a^{2}+r_{0}^{2}\right)^{\frac{D+3}{2}}T_{0}}{2Da^{4}+\left(D^{2}-3D+6\right)a^{2}r_{0}^{2}-(D-3)r_{0}^{4}}\,.$ (109)
We see that for low temperature, or equivalently smaller $r_{0}$, $C_{J}$ is positive. The thermodynamics is stable when $r_{0}$ lies in the range
$\frac{2a^{2}}{D-3}\equiv r_{\dagger}^{2}\leq r_{0}^{2}\leq r_{\sharp}^{2}\equiv\frac{D^{2}-3D+6+\sqrt{(D-2)(D-1)\left(D^{2}-3D+18\right)}}{2(D-3)}a^{2}\,.$ (110)
Here $r_{\dagger}$ is the radius parameter of the extremal black hole $(T_{0}=0)$. Beyond $r_{\sharp}$, the capacity is negative and the thermodynamics becomes unstable. On the other hand, it follows from (105), the sign of $\Delta S/\alpha$ in the full ranges of thermodynamic stability is given by
${\frac{\Delta S}{\alpha}}<0:$ $\quad\frac{2a^{2}}{D-3}\leq r_{0}^{2}<r_{\ddagger}^{2}\equiv\frac{(2D-3)+\sqrt{(D-1)(3D-5)}}{(D-2)^{2}}a^{2}<r_{\sharp}^{2}\,,$ (111)
${\frac{\Delta S}{\alpha}}>0:$ $\quad r_{\ddagger}^{2}<r_{0}^{2}\leq r_{\sharp}^{2}\,.$ (112)
Thus we see that $\Delta S$ cannot be always positive in the stability region regardless the sign choice of the coupling constant $\alpha$.
In even $D=2N+2$ dimensions with all equal angular momenta, the story is similar. We have the specific heat capacity
$C_{J}=\frac{\pi(D-2)\Sigma_{D-2}r_{0}\left(a^{2}+r_{0}^{2}\right)^{\frac{D}{2}+1}T_{0}}{a^{4}(D-1)+\left(D^{2}-4D+6\right)a^{2}r_{0}^{2}-(D-3)r_{0}^{4}}\,,$ (113)
which implies that the the thermodynamic is stable in the following parameter region:
$\frac{a^{2}}{D-3}\equiv r_{\dagger}^{2}\leq r_{0}^{2}\leq r_{\sharp}^{2}\equiv\frac{6-4D+D^{2}+(-2+D)\sqrt{12-4D+D^{2}}}{2(D-3)}a^{2}\,.$ (114)
It follows from (106) that $\Delta S$ can change its sign when $r_{0}$ crosses over $r_{\ddagger}$, with
$r_{\dagger}^{2}<r_{\ddagger}^{2}\equiv\frac{2a^{2}}{D-2}<r_{\sharp}^{2}\,.$ (115)
We thus find counterexamples of [19], although it can be argued that it is much more difficult to fix the angular momenta than the electric charges in a thermodynamic process. A counterexample to [19] with varying angular momenta was found in the higher-derivative correction in supergravity [26].
@@ 6 Kerr black hole under cubic curvature correction
It is worth pointing out again that the specific heat capacity is
$C_{J}=\frac{8\pi^{2}r_{0}\left(a^{2}+r_{0}^{2}\right){}^{3}}{3a^{4}+6a^{2}r_{0}^{2}-r_{0}^{4}}\,.$ (120)
We therefore have stable thermodynamics for the fixed angular momentum in the following parameter region.
$a^{2}\leq r_{0}^{2}\leq r_{\sharp}^{2}\equiv\left(3+2\sqrt{3}\right)a^{2}\,.$ (121)
However $\Delta S$ is not positive definite in the stable region. $\Delta S$ charges its sign when $r_{0}$ cross over $r_{\ddagger}\sim 2.077a$.
@@ 7 Divergence is a perturbative artefact: an explicit example
While we do not have exact solutions of rotating black holes in Einstein gravity extended with quadratic or more general higher-order derivative terms, we expect that the divergence we explored in this paper is an artefact of perturbation, indicating a breakdown only of the perturbative approach. The divergence may not necessarily exist in the full nonlinear level.
@@ 8 Conclusions
In this paper, we began a review of Myers-Perry rotating black holes in diverse dimensions. We presented a simple analytic proof that the rotating black holes in even dimensions are all thermodynamically unstable. Specifically, we showed that the trace of the corresponding thermodynamic metric has an elegant expression (21) and it is manifestly non-positive. However, for odd dimensions, we have to employ numerical analysis to prove the instability. We also found that the specific heat capacity of fixed angular momenta was positive at low temperature, including the extremal limit. This implies that the thermodynamics could be stable for the reduced system where all the angular momenta are fixed to some appropriate specific values and hence extracted from the thermodynamic process. We also found that in odd dimensions when one and only one angular momentum vanishes, there existed an intriguing decoupling limit of extremal solutions whose geometry (16) has an unusual AdS3 embedding.
@@
In addition to the divergence of the $\alpha$-order correction, there are points in the parameter space that the correction vanishes. What is intriguing is that these zeros occur at the same points for the corrections to different thermodynamic variables in different ensembles. For example, $\Delta J$ in grand canonical ensemble vanishes at the exact same point as $\Delta\Omega$ in canonical ensemble. This can be seen from the numerator of $\Delta J$ in eqns (3.2.1), (3.2.1), (41), (3.3), and compare them to the corresponding $\Delta\Omega$ in eqns (70), (74), (79), (86). The similar relation exists between $\Delta M$ in canonical ensemble and $\Delta T$ in system with fixed mass and angular momenta, as we can see from eqns (70), (74), (79), (86) and (5.1), (5.1), (5.1), (5.1). These phenomena can be understood from the identities of differential equations, namely
@@
We also examined the issue of the sign of $\Delta S$ under the quadratic curvature extension. Since there is only one nontrivial coupling constant $\alpha$, this would provide a strong condition on the sign choice of the $\alpha$. However, for the reduce stable thermodynamic system where all the angular momenta are taken to be some specific values, we found that $\Delta S$ could change sign within the stable parameter region, therefore providing counterexamples of [19].